OpenStudy (anonymous):

Can some one help me with these two problems? 1. sqrt(64m^4r^12) 2. sqrt(200x^8y^15)

OpenStudy (anonymous):

\[\sqrt{64m ^{4}r ^{12}}\] What times itself will give you 64?

OpenStudy (anonymous):

Perfect squares include 1*1=1 2*2=4 3*3=9 4*4=16 5*5=25 and so on

OpenStudy (anonymous):

I know it's 8m^2 but what about the r^12? Does it turn into 3 & 4 or 3 6 and 2 ? I got confused, and I don't know what to do what to do after.

OpenStudy (anonymous):

3 and 6*

OpenStudy (anonymous):

take the 12 and divide by 2 you get 8m^2r^6

OpenStudy (anonymous):

finding the square roots of exponents under the sign, divide by 2

OpenStudy (anonymous):

So I don't have to divide the 6?

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

sqrt(200x^8y^15) 200 isn't a perfect square--but you can take 100*2 and find the sq rt of 100

OpenStudy (anonymous):

\[\sqrt{200x ^{8}y ^{15}}\] \[10x ^{4}y ^{7}\sqrt{2y}\]